OpenStudy (anonymous):

Find all solutions of sin^3(x) +s sin^2(x) - sinx -1 = 0 on the interval [0, 2π]

OpenStudy (anonymous):

\[\sin^3x + \sin^2x -sinx -1 =0\] =\[(\sin^2x-1)(sinx+1)\] \[\sin^2x=1 | sinx = -1\] ...................................... \[sinx = -1, x=3 \pi/2\] ...................................... \[\sin^2x = 1 , sinx = \pm1\] \[x=\pi/2\] ....................................... Thus x= \[\pm \pi/2\]

OpenStudy (anonymous):

Thank you!